how many gallons of 1/2% milk [1/2% fat]should be mixed with 24 gallons of 2% milk [2% fat to produce a mixture that is 1% fat
Answer by richwmiller(17219)

We have 24 gal of 2% milk fat
We want to add some 0.5% milk fat to make 1% milk fat
0.005*x+0.02*24=0.01(24+x)
0.005*x+0.48=0.24+0.01x
0.005x-0.01x=0.24-0.48
-0.005x=-0.24
x=48.0 gal at 0.5% (0.24) is added to 24 gal at 2%(0.48) to make 1% of 72 gal(0.72)
check
0.005*48+0.02*24=0.01(24+48)
0.24+0.48=0.01(72)
0.72=0.72
ok
codemixadd